# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 3.1. Additivity
The question of the additivity of strong homology has, over the past 35 years, attracted a substantial literature; main results therein include the following:
@@
    The continuum hypothesis implies that strong homology is not additive [54].
@@
We turn now to our central examples. For $n>0$ let $B^{n}$ denote an $n$-dimensional open ball, and let $S^{n}$, as usual, denote an $n$-dimensional sphere. For any cardinal $\lambda>0$ let $\mathbf{Y}^{n,\lambda}$ denote the inverse system $(Y^{n,\lambda}_{a},p_{ab},[\lambda]^{<\omega}\backslash\varnothing)$ in which for all nonempty $a\subseteq b$ in $[\lambda]^{<\omega}$ the space $Y^{n,\lambda}_{a}$ is $\bigvee_{a}S^{n}$ and the bonding map $p_{ab}:Y^{n,\lambda}_{b}\to Y^{n,\lambda}_{a}$ is the identity on $\bigvee_{a}S^{n}$, sending all other points in $\bigvee_{b}Y^{n,\lambda}$ to the wedge’s basepoint.
@@
The second claim follows from the first together with [56, Thm. 6.20]. For the first, observe simply that $Y^{n,\lambda}$ is compact, that the $\alpha^{\mathrm{th}}$ summand of $\coprod_{\lambda}B^{n}$ naturally identifies with the complement of the basepoint in $Y^{n,\lambda}_{\{\alpha\}}$, and that the complement of the union of these $B^{n}$-images in $Y^{n,\lambda}$ is exactly one point. ∎
@@
The last assertion follows from the first simply by letting $n$ range above 00 and letting $q$ range between them. For the first, observe that by Lemma 3.2 the arguments of [54, §5-6] fully generalize to arbitrary cardinalities $\kappa$ and $\lambda$, with only the partial exception of the (implicit) invocation of the Mittag-Leffler principle at the top of page 736. In our context, the key point at that step is that for every $q\geq 0$ the derived limits $\mathrm{lim}^{s}$ of the system of singular homology groups $\mathrm{H}_{q}(Y^{n,\lambda}_{b})$ $(b\in[\lambda]^{<\omega})$ associated to $\mathbf{Y}^{n,\lambda}$ vanish in all degrees $s>0$. When $q\neq n$, though, this is clear, since these groups are either constantly 00 or $\mathbb{Z}$, while if $q=n$, this system is $\mathbf{A}_{1,\lambda}$, whereupon the assertion follows immediately from item (1) of Theorem 2.10. ∎
@@
Our counterexample is a countable sum of copies of $Y^{n,\aleph_{1}}$. By Theorem 3.3, whenever $p=n-1$, the additivity of strong homology over this sum would require that $\mathrm{lim}^{1}\,\mathbf{A}_{\aleph_{0},\aleph_{1}}=0$, but we know from Theorem 2.10 that this limit is nonzero. ∎
As noted, $Y^{n,\omega_{1}}$ is a significantly simpler counterexample to additivity than that appearing in [67] (which had derived, in turn, from [55]); it is also, in contrast to that counterexample, compact. See our conclusion for the question of whether metrizable ZFC counterexamples to additivity exist. There exists no locally compact separable metric ZFC counterexample to the additivity of strong homology since, by [6], the only obstructions to additivity on this class of spaces are any nonvanishing derived limits $\mathrm{lim}^{n}$ $(n>0)$ of a family of inverse systems which sufficiently resemble $\mathbf{A}_{\aleph_{0},\aleph_{0}}$ so that their higher limits all vanish for the same reasons that the latter’s do in the model of [16] (and likewise in the model of [14], by [5]). When $\mathrm{lim}^{1}\,\mathbf{A}_{\aleph_{0},\aleph_{0}}\neq 0$, on the other hand, strong homology fails to be additive even on closed subspaces of $\mathbb{R}^{2}$, as witnessed by a countable sum of earrings $Y^{1,\aleph_{0}}$, and it was the derivation of this scenario from the continuum hypothesis which formed the main result of [54].
These examples, like those of [67], may be regarded as materializing a more fundamental failure of additivity, namely one at the level of the derived limit functors. More particularly, [67, Theorem 1] is also immediate within our framework, this time from Theorem 2.10.
@@
Note in contrast that the functor $\mathrm{lim}^{0}=\mathrm{lim}:\mathsf{Pro(Ab)}\to\mathsf{Ab}$ is additive, as follows from the observation beginning the proof below. [^3] The additivity of Čech homology is argued in [54, Theorem 9]; the proof applies in its essentials to the functor $\mathrm{lim}$ as well, but readers may find it at least as efficient to supply the argument themselves.
@@
The essential observation is that the sum, in a pro-category, of a set of inverse systems is represented by a system indexed by the product of their index sets (with terms the naturally associated sum of objects; see [67, pp. 502-503] for details). In particular, $\mathbf{A}_{\kappa,\lambda}=\coprod_{\kappa}\mathbf{A}_{1,\lambda}$ for all cardinals $\kappa$ and $\lambda$, so the inequality $\mathrm{lim}^{1}\,\mathbf{A}_{\aleph_{0},\aleph_{1}}\neq 0=\bigoplus_{\omega}\mathrm{lim}^{1}\,\mathbf{A}_{1,\aleph_{1}}$ witnesses the $n=1$ instance of the corollary. For the $n=2$ case, let $\mathbf{C}_{\kappa,\lambda}=(C_{f},q_{fg},{{}^{\kappa}}([\lambda]^{<\omega})$ and $\mathbf{D}_{\kappa,\lambda}=(D_{f},r_{fg},{{}^{\kappa}}([\lambda]^{<\omega})$, where
@@
An oddity of these results, of course, is their restriction to the degrees $n\leq 2$. By Goblot’s Theorem, results for higher $n$ will be, just as in Section 2, at least in part about the combinatorics of $\aleph_{n-1}$, and this indeed is much of the interest of Question 5.6 below. In particular, the question of whether the functor $\mathrm{lim}^{3}:\mathsf{pro\text{-}Ab}\to\mathsf{Ab}$ is consistently additive appears to be closely related to the question of whether the group $\mathrm{H}^{2}(\omega_{2};\mathbb{Z})$ can vanish in any model of the ZFC axioms. [^4] See [12, Question 9.3] and [14, Question 7.7] and [13, Question 11.2] for general formulations of this question, which is one of the most important in this research area.
Of comparable interest is the question on what classes $\mathcal{C}$ may the functors $\mathrm{lim}^{n}$ be additive? After all, the ZFC obstructions recorded in Theorem 2.10 and Corollary 3.5 only arise for systems of “height” $\lambda\geq\aleph_{1}$; it is natural, in consequence, to wonder if $\mathrm{lim}^{n}$ may nevertheless be additive for restricted or unrestricted sums of towers (i.e., countable inverse sequences) of abelian groups, and it’s indeed in these terms that the theorem [5, Thm. 1.2] referenced in Section 2.4 is framed. In fact (as noted in [5, §7]), this result coupled with the argument of [11, Thm. 5.1] implies the consistency of $\mathrm{lim}^{1}\,\mathbf{A}_{\kappa,\aleph_{0}}[H]=0$ for all $\kappa$ and abelian groups $H$, and the question of whether this can hold for higher degrees as well appears to be a deep one.
@@ 3.2. Strong homology
Works on the additivity problem have, following [54] and [56], tended to emphasize the axiomatic and, to a lesser degree, computational features of strong homology, somewhat eliding more conceptual formulations in the process (see [6, §1] for a recent example). The cumulative effect has been a distorted picture, one in which strong homology figures mainly (1) as a sort of compromise-formation between the requirements of a homology functor and the unruly realities of geometric topology (represented by a standard roster of motivating examples of “bad local behavior”: the Warsaw circle, Hawaiian earrings, solenoids, and so on), or (2) as an elaborate, if not ad hoc, exactness-recovering corrective to Čech homology, or (3) as both. There is an element of truth to each of these images, but taken alone, they are seriously misleading as well, insofar as strong homology figures within them as essentially isolated from the more structural and homotopy theoretic emphases of the contemporary mainstream of algebraic topology. The reality, though, is that those emphases and strong homology grew up together in large part out of the arrival, through the 1970s, of model category, homotopy (co)limit, and homotopy coherent technologies [18, 20, 28], and it is in these terms, at least as much as (1) and (2) above, that strong homology should be understood; that strong shape, the background homotopy theory of strong homology, is a conspicuously $\infty$-categorical entity is just one manifestation of this heritage (see, e.g., [50, §7.1.6] and [40]). We briefly sketch a more balanced, but nevertheless classical, account below.
@@
Surely the best-known general invariant of a topological space $X$ is its homotopy type; put differently, the most prominent relation refining the identity relation on topological spaces is that of homotopy equivalence. For all its virtues, this relation is, from some perspectives, weak; natural extensions of this relation determine further invariants of a topological space $X$, such as its strong shape and its shape. [^5] We note in passing that these extensions are, in spirit, orthogonal to that of weak homotopy equivalence: while the latter identifies the Warsaw circle, for example, with a point, the former identify it with a circle. These more intuitive identifications are the source both of the name shape and of much initial interest in its theory, concerns emphasized in [27] and recapitulated (alongside many fundamental techniques) in the more contemporary field of persistent homology. More precisely, we have a sequence of categories $\mathsf{Ho(Top)}\to\mathsf{sSh(Top)}\to\mathsf{Sh(Top)}$ corresponding to the aforementioned invariants, respectively, together with commuting functors from $\mathsf{Top}$ to each of them. The middle of these admits multiple incarnations, or models (within the coherent homotopy category $\mathsf{CH}(\mathsf{Pro}(\mathsf{Top}))$, for example; see [56, pp. 90, 2–3, and Ch. 1]); most classical among these is that of [33], whereby the above diagram takes the form
[…]
@@
with each of the latter two terms naturally construed as categories of systems of approximations to a topological space $X$ [^6] To be clear, the homotopy category of the first and third terms is, in the terminology of [61, p. 395], the “naive” one, corresponding to the Hurewicz model structure on the category of topological spaces; that of the middle term derives from either of the classes of weak equivalences discussed in [66]. For the functor $F$, see [56, §8]; on objects, $G$ factors through the natural map $\textsf{Pro}(\mathsf{Top})\to\mathsf{Pro}(\mathsf{Ho(Top)})$. Such systems are only informative, of course, when they take values in some restricted class (or dense subcategory) of “nice” topological spaces, the categories $\mathsf{Pol}$ of polyhedra or of almost neighborhood retracts (ANRs) being standard historical choices. And indeed, the Vietoris nerve and Čech nerve constructions on the objects of $\mathsf{Top}$ may each naturally be viewed as functors to the Ho(Pro(Pol)) and Pro(Ho(Pol)) subcategories, respectively, of the incarnations of $\mathsf{sSh(Top)}$ and $\mathsf{Sh(Top)}$ in equation (12) above (see [41]). As is well known, within the Čech construction, the plurality of maps witnessing the refinement relation between two open coverings of a space induce a single homotopy class of maps between the associated polyhedra, and it is for this reason that the shape functor takes values in the pro-category of $\mathsf{Ho(Top)}$.
@@
The further passage to a homology theory then depends on the assignment of (homotopy classes of) chain complexes to the objects of the terms of equation (12). The canonical choice in the case of $\mathsf{Ho(Top)}$ is, of course, the singular chain functor $X\mapsto C^{\mathrm{sing}}_{\bullet}(X)$, the crucial point here being that $C^{\mathrm{sing}}_{\bullet}$ maps homotopy equivalent spaces to homotopy equivalent chain complexes; the homology groups of $C^{\mathrm{sing}}_{\bullet}(X)$ are the singular homology groups of $X$. On the other hand, applying $C^{\mathrm{sing}}_{\bullet}$ to the $\mathsf{Top}$ and $\mathsf{Ho(Top)}$ terms of the two subsequent terms of line (12) yields objects in the homotopy category of the pro-category of chain complexes and the pro-category of the homotopy category of chain complexes, respectively. The natural means for consolidating these systems into a single graded abelian group are the homotopy inverse limit and the inverse limit, respectively — although, limits not in general existing in homotopy categories, it’s in the latter case standard to interchange steps, i.e., to take the limit of the homology groups of the chain complexes of the elements of the Čech nerve of $X$; the resultant groups are the Čech homology groups of $X$ [^7] This parallels, of course, the standard construction of Čech cohomology groups; that the latter are both shape and strong shape invariant is essentially immediate.
@@
where $\mathrm{H}_{q}$ denotes singular homology, $S:\mathsf{Top}\to\mathsf{Sh(Top)}$ and $sS:\mathsf{Top}\to\mathsf{sSh(Top)}$ denote the shape (see [57, I.4.2, Thm. 3]) and strong shape functors, respectively, and $h_{q}$ denotes the $q^{\mathrm{th}}$ homology group of a chain complex. Cordier (who characterizes the $\mathrm{holim}$ and $\mathrm{lim}^{s}$ functors as complementary responses to the “deficiencies” of $\mathrm{lim}$ [29, p. 35]) derives the existence of a homotopy limit in the category of non-negative chain complexes of abelian groups from the Dold-Kan correspondence together with its existence in the category of simplicial abelian groups; he shows moreover that it takes the concrete form of precisely that total complex figuring so prominently in Chapter 4 of [56].
@@
  • Generalized framings of the strong shape functor are of a wider significance than we have so far suggested; see in particular Hoyois’s Higher Galois theory [44, Def. 2.3, Rmk. 2.13], where this functor, ranging over $\infty$-topoi and denoted $\Pi_{\infty}$, carries the name fundamental pro-$\infty$-groupoid; shape functors arise naturally in both the condensed and pyknotic settings as well, and are areas of active research. [^8] See forthcoming work by Mair; see also [8, II.4] for further references.
  • Recall that Eilenberg and Steenrod showed, via a geometric realization of a simple inverse sequence of groups possessing a nonvanishing $\mathrm{lim}^{1}$ “that the Čech ‘homology theory’ with integer coefficients is not exact,” or more precisely that no integral homology theory on the category of compact pairs is simultaneously continuous (meaning that it commutes with inverse limits) and exact [34, p. 265]. Consider then what is sometimes framed (in a sense made precise by a Milnor short exact sequence [56, Thm. 21.9]) as rectifying the non-exactness of $\check{\mathrm{H}}_{\bullet}$, namely the strong homology functor $\overline{\mathrm{H}}_{\bullet}$. On the category of compact metric spaces, $\overline{\mathrm{H}}_{\bullet}$ is fully axiomatized by the Eilenberg-Steenrod axioms together with a relative homeomorphism and cluster axiom [64]; it is in consequence thereon equivalent to the homology theories of Steenrod [78], Borel-Moore [19], Sitnikov, and others [59]. Where these theories may differ is in their extension to the broader category of locally compact spaces; of the central question of whether the Steenrod-Sitnikov theory embodying the most naive strategy of extension (direct limits) is strong shape invariant, Sklyarenko writes that it “will likely be a test not only for Steenrod-Sitnikov homology, but also for strong shape theory itself” ([76], quoted in [62]). It is reasonable to wonder whether materializations like $\coprod_{\omega}Y^{n,\aleph_{1}}$ of nonvanishing derived limits may, much as in the classical Čech case, play some role in the answer.

### v2

@@ 3.1. Additivity
The question of the additivity of strong homology has, over the past 35 years, attracted a substantial literature; in chronological order, main results therein include the following:
@@
    The continuum hypothesis implies that strong homology is not additive, not even on the category of closed subsets of $\mathbb{R}^{2}$ [57].
@@
[…]
@@
The second claim follows from the first together with [59, Thm. 6.20]. For the first, observe simply that $Y^{n,\lambda}$ is compact, that the $\alpha^{\mathrm{th}}$ summand of $\coprod_{\lambda}B^{n}$ naturally identifies with the complement of the basepoint in $Y^{n,\lambda}_{\{\alpha\}}$, and that the complement of the union of these $B^{n}$-images in $Y^{n,\lambda}$ is exactly one point. ∎
@@
The last assertion follows from the first simply by letting $n$ range above 0 and letting $q$ range between them. For the first, observe that by Lemma 3.2 the arguments of [57, §5-6] fully generalize to arbitrary cardinalities $\kappa$ and $\lambda$, with only the partial exception of the (implicit) invocation of the Mittag-Leffler principle at the top of page 736. In our context, the key point at that step is that for every $q\geq 0$ the derived limits $\mathrm{lim}^{s}$ of the system of singular homology groups $\mathrm{H}_{q}(Y^{n,\lambda}_{b})$ $(b\in[\lambda]^{<\omega})$ associated to $\mathbf{Y}^{n,\lambda}$ vanish in all degrees $s>0$. When $q\neq n$, though, this is clear, since these groups are either constantly 0 or $\mathbb{Z}$, while if $q=n$, this system is $\mathbf{A}_{1,\lambda}$, whereupon the assertion follows immediately from item (1) of Theorem 2.10. ∎
@@
Our counterexample is a countable sum of copies of $Y^{n,\aleph_{1}}$. By Theorem 3.3, whenever p=n1p=n-1, the additivity of strong homology over this sum would require that $\mathrm{lim}^{1}\,\mathbf{A}_{\aleph_{0},\aleph_{1}}=0$, but we know from Theorem 2.10 that this limit is nonzero. ∎
As noted, $Y^{n,\aleph_{1}}$ is a significantly simpler counterexample to additivity than that appearing in [70] (which had derived, in turn, from [58]); it is also, in contrast to that counterexample, compact. See our conclusion for the question of whether metrizable ZFC counterexamples to additivity exist. There exists no locally compact separable metric ZFC counterexample to the additivity of strong homology since, by [7], the only obstructions to additivity on this class of spaces are any nonvanishing derived limits $\mathrm{lim}^{n}$ $(n>0)$ of a family of inverse systems which sufficiently resemble $\mathbf{A}_{\aleph_{0},\aleph_{0}}$ so that their higher limits all vanish for the same reasons that the latter’s do in the model of [18] (and likewise in the model of [16], by [5]). When $\mathrm{lim}^{1}\,\mathbf{A}_{\aleph_{0},\aleph_{0}}\neq 0$, on the other hand, strong homology fails to be additive even on closed subspaces of $\mathbb{R}^{2}$, as witnessed by a countable sum of earrings $Y^{1,\aleph_{0}}$, and it was the derivation of this scenario from the continuum hypothesis which formed the main result of [57].
These examples, like those of [70], may be regarded as materializing a more fundamental failure of additivity, namely one at the level of the derived limit functors. More particularly, [70, Theorem 1] is also immediate within our framework, this time from Theorem 2.10.
@@
Note in contrast that the functor $\mathrm{lim}^{0}=\mathrm{lim}:\mathsf{Pro(Ab)}\to\mathsf{Ab}$ is additive, as follows from the observation beginning the proof below. [^3] The additivity of Čech homology is argued in [57, Theorem 9]; the proof applies in its essentials to the functor $\mathrm{lim}$ as well, but readers may find it at least as efficient to supply the argument themselves.
@@
[…]
@@
An oddity of these results, of course, is their restriction to the degrees $n\leq 2$. By Goblot’s Theorem, results for higher $n$ will be, just as in Section 2, at least in part about the combinatorics of $\aleph_{n-1}$, and this indeed is much of the interest of Question 5.6 below. In particular, the question of whether the functor $\mathrm{lim}^{3}:\mathsf{Pro(Ab)}\to\mathsf{Ab}$ is consistently additive appears to be closely related to the question of whether the group $\mathrm{H}^{2}(\omega_{2};\mathbb{Z})$ can vanish in any model of the ZFC axioms. [^4] See [13, Question 9.3] and [16, Question 7.7] and [14, Question 11.2] for general formulations of this question, which is one of the most important in this research area.
Of comparable interest is the question on what classes $\mathcal{C}$ may the functors $\mathrm{lim}^{n}$ be additive? After all, the ZFC obstructions recorded in Theorem 2.10 and Corollary 3.5 only arise for systems of “height” $\lambda\geq\aleph_{1}$; it is natural, in consequence, to wonder if $\mathrm{lim}^{n}$ may nevertheless be additive for restricted or unrestricted sums of towers (i.e., countable inverse sequences) of abelian groups, and it’s indeed in these terms that the theorem [5, Thm. 1.2] referenced in Section 2.4 is framed. In fact (as noted in [5, §7]), this result coupled with the argument of [12, Thm. 5.1] implies the consistency of $\mathrm{lim}^{1}\,\mathbf{A}_{\kappa,\aleph_{0}}[H]=0$ for all $\kappa$ and abelian groups $H$, and the question of whether similar implications hold in higher degrees appears to be a deep one; for recent developments, see [6, 15].
@@ 3.2. Strong homology
Works on the additivity problem have, following [57] and [59], tended to emphasize the axiomatic and, to a lesser degree, computational features of strong homology, somewhat eliding more conceptual formulations in the process (see [7, §1] for a recent example). The cumulative effect has been a distorted picture, one in which strong homology figures mainly (1) as a sort of compromise-formation between the requirements of a homology functor and the unruly realities of geometric topology (represented by a standard roster of motivating examples of “bad local behavior”: the Warsaw circle, Hawaiian earrings, solenoids, and so on), or (2) as an elaborate, if not ad hoc, exactness-recovering corrective to Čech homology, or (3) as both. There is an element of truth to each of these images, but taken alone, they are seriously misleading as well, insofar as strong homology figures within them as essentially isolated from the more structural and homotopy theoretic emphases of the contemporary mainstream of algebraic topology. The reality, though, is that those emphases and strong homology grew up together in large part out of the arrival, through the 1970s, of model category, homotopy (co)limit, and homotopy coherent technologies [20, 22, 30], and it is in these terms, at least as much as (1) and (2) above, that strong homology should be understood; that strong shape, the background homotopy theory of strong homology, is a conspicuously $\infty$-categorical entity is just one manifestation of this heritage (see, e.g., [53, §7.1.6] and [43]). We briefly sketch a more balanced, but nevertheless classical, account below.
@@
Surely the best-known general invariant of a topological space $X$ is its homotopy type; put differently, the most prominent relation refining the identity relation on topological spaces is that of homotopy equivalence. For all its virtues, this relation is, from some perspectives, weak; natural extensions of this relation determine further invariants of a topological space $X$, such as its strong shape and its shape. [^5] We note in passing that these extensions are, in spirit, orthogonal to that of weak homotopy equivalence: while the latter identifies the Warsaw circle, for example, with a point, the former identify it with a circle. These more intuitive identifications are the source both of the name shape and of much initial interest in its theory, concerns emphasized in [29] and recapitulated (alongside many fundamental techniques) in the more contemporary field of persistent homology. More precisely, we have a sequence of categories $\mathsf{Ho(Top)}\to\mathsf{sSh(Top)}\to\mathsf{Sh(Top)}$ corresponding to the aforementioned invariants, respectively, together with commuting functors from $\mathsf{Top}$ to each of them. The middle of these admits multiple incarnations, or models (within the coherent homotopy category $\mathsf{CH}(\mathsf{Pro}(\mathsf{Top}))$, for example; see [59, pp. 90, 2–3, and Ch. 1]); most classical among these is that of [35], whereby the above diagram takes the form
(12) $\mathsf{Ho(Top)}\xrightarrow{F}\textsf{Ho(Pro(Top))}\xrightarrow{G}\textsf{Pro(Ho(Top))},$
@@
with each of the latter two terms naturally construed as categories of systems of approximations to a topological space XX. [^6] To be clear, the homotopy category of the first and third terms is, in the terminology of [64, p. 395], the “naive” one, corresponding to the Hurewicz model structure on the category of topological spaces; that of the middle term derives from either of the classes of weak equivalences discussed in [69]. For the functor $F$, see [59, §8]; on objects, $G$ factors through the natural map $\textsf{Pro}(\mathsf{Top})\to\mathsf{Pro}(\mathsf{Ho(Top)})$. Such systems are only informative, of course, when they take values in some restricted class (or dense subcategory) of “nice” topological spaces, the categories $\mathsf{Pol}$ of polyhedra or of absolute neighborhood retracts (ANRs) being standard historical choices. And indeed, the Vietoris nerve and Čech nerve constructions on the objects of $\mathsf{Top}$ may each naturally be viewed as functors to the Ho(Pro(Pol)) and Pro(Ho(Pol)) subcategories, respectively, of the incarnations of $\mathsf{sSh(Top)}$ and $\mathsf{Sh(Top)}$ in equation (12) above (see [44]). As is well known, within the Čech construction, the plurality of maps witnessing the refinement relation between two open coverings of a space induce a single homotopy class of maps between the associated polyhedra, and it is for this reason that the shape functor takes values in the pro-category of $\mathsf{Ho(Top)}$.
@@
The further passage to a homology theory then depends on the assignment of (homotopy classes of) chain complexes to the objects of the terms of equation (12). The canonical choice in the case of $\mathsf{Ho(Top)}$ is, of course, the singular chain functor $X\mapsto C^{\mathrm{sing}}_{\bullet}(X)$, the crucial point here being that $C^{\mathrm{sing}}_{\bullet}$ maps homotopy equivalent spaces to homotopy equivalent chain complexes; the homology groups of $C^{\mathrm{sing}}_{\bullet}(X)$ are the singular homology groups of $X$. On the other hand, applying $C^{\mathrm{sing}}_{\bullet}$ to the $\mathsf{Top}$ and $\mathsf{Ho(Top)}$ terms of the two subsequent terms of line (12) yields objects in the homotopy category of the pro-category of chain complexes and the pro-category of the homotopy category of chain complexes, respectively. The natural means for consolidating these systems into a single graded abelian group are the homotopy inverse limit and the inverse limit, respectively — although, limits not in general existing in homotopy categories, it’s in the latter case standard to interchange steps, i.e., to take the limit of the homology groups of the chain complexes of the elements of the Čech nerve of $X$; the resultant groups are the Čech homology groups of XX. [^7] This parallels, of course, the standard construction of Čech cohomology groups; that the latter are both shape and strong shape invariant is essentially immediate.
@@
where $\mathrm{H}_{q}$ denotes singular homology, $S:\mathsf{Top}\to\mathsf{Sh(Top)}$ and $sS:\mathsf{Top}\to\mathsf{sSh(Top)}$ denote the shape (see [60, I.4.2, Thm. 3]) and strong shape functors, respectively, and $h_{q}$ denotes the $q^{\mathrm{th}}$ homology group of a chain complex. Cordier (who characterizes the $\mathrm{holim}$ and $\mathrm{lim}^{s}$ functors as complementary responses to the “deficiencies” of $\mathrm{lim}$ [31, p. 35]) derives the existence of a homotopy limit in the category of non-negative chain complexes of abelian groups from the Dold-Kan correspondence together with its existence in the category of simplicial abelian groups; he shows moreover that it takes the concrete form of precisely that total complex figuring so prominently in Chapter 4 of [59].
@@
  • Generalized framings of the strong shape functor are of a wider significance than we have so far suggested; see in particular Hoyois’s Higher Galois theory [47, Def. 2.3, Rmk. 2.13], where this functor, ranging over $\infty$-topoi and denoted $\Pi_{\infty}$, carries the name fundamental pro-$\infty$-groupoid; shape functors arise naturally in both the condensed and pyknotic settings as well, and are areas of active research. [^8] See forthcoming work by Mair; see also [9, II.4] for further references.
  • Recall that Eilenberg and Steenrod showed, via a geometric realization of a simple inverse sequence of groups possessing a nonvanishing $\mathrm{lim}^{1}$ “that the Čech ‘homology theory’ with integer coefficients is not exact,” or more precisely that no integral homology theory on the category of compact pairs is simultaneously continuous (meaning that it commutes with inverse limits) and exact [36, p. 265]. Consider then what is sometimes framed (in a sense made precise by a Milnor short exact sequence [59, Thm. 21.9]) as rectifying the non-exactness of $\check{\mathrm{H}}_{\bullet}$, namely the strong homology functor $\overline{\mathrm{H}}_{\bullet}$. On the category of compact metric spaces, $\overline{\mathrm{H}}_{\bullet}$ is fully axiomatized by the Eilenberg-Steenrod axioms together with a relative homeomorphism and cluster axiom [67]; it is in consequence thereon equivalent to the homology theories of Steenrod [81], Borel-Moore [21], Sitnikov, and others [62]. Where these theories may differ is in their extension to the broader category of locally compact spaces; of the central question of whether the Steenrod-Sitnikov theory embodying the most naive strategy of extension (direct limits) is strong shape invariant, Sklyarenko writes that it “will likely be a test not only for Steenrod-Sitnikov homology, but also for strong shape theory itself” ([79], quoted in [65]). It is reasonable to wonder whether materializations like $\coprod_{\omega}Y^{n,\aleph_{1}}$ of nonvanishing derived limits may, much as in the classical Čech case, play some role in the answer.